A television set that originally cost $2,340 was reduced in price by successive discounts of 20%, 25% and 10%. What was the equivalent single rate of discount?
Answer by richwmiller(17219)

2340*.80*.75*.90=1263.60
1263.60/2340=0.54
54% of original price which is a discount of 46%
.80*.75*.90=.54
.46*2340=1076.40
check
2340-1076.40=1263.60
ok